Solve for the confidence interval estimate of the population proportion.


Mikay conducted poll survey in which 520 of 1200 randomly selected voters indicated their preference for a certain candidate. What is the true population proportion of voters who prefer the candidate using 95% confidence interval?


1
Expert's answer
2021-05-07T12:01:54-0400

We have given that,


"p = \dfrac{520}{1200} = 0.43"


Confidence interval can be given as,


"CI = p \pm z \sqrt{\dfrac{p(1-p)}{n}}"


"CI = 0.43 \pm 1.96 \sqrt{\dfrac{0.43 \times 0.57}{1200}}"


"CI = 0.43 \pm 0.028"